OpenStudy (anonymous):

Is the expression x^3•x^3•x^3 equivalent to x^3•3•3? Why or why not? Explain your reasoning.

zepdrix (zepdrix):

Hey Steve, It's a little difficult to understand where your 3's are placed. Does the problem look like this? \(\large\rm x^3\cdot x^3\cdot x^3\) and \(\large\rm x^{3\cdot3\cdot3}\)

zepdrix (zepdrix):

Here is a handy exponent rule to remember: \(\large\rm x^a\cdot x^b\quad=\quad x^{a+b}\) The exponents get `added together`, not multiplied.

OpenStudy (anonymous):

yes it looks like what you said

OpenStudy (anonymous):

i know what the answer is for \[x ^{3}x ^{3}x ^{3}\] but i dont know how to find the answer to the second equation

OpenStudy (anonymous):

is it just x^9?

zepdrix (zepdrix):

Using our exponent rule we can determine that \(\large\rm x^3\cdot x^3\cdot x^3=x^{3+3+3}\) good good, x^9. which is NOT the same as \(\large\rm x^{3\cdot3\cdot3}\) right?

OpenStudy (anonymous):

yes ok thank you very much

zepdrix (zepdrix):

3*3*3=27 np